OpenStudy (anonymous):

can someone help me solve and understand this? will fan and give medal!:) What is the sum of the finite arithmetic series? 26 + 29 + 32 + 35 + 38 + 41 + 44 (1 point) 219 245 –193 201

OpenStudy (anonymous):

245

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

ADDED them checked my answer with mathway

OpenStudy (anonymous):

(26+44)+(29+41)+(32+38)+35 =70*3+35 =?

OpenStudy (anonymous):

or \[Sn=\frac{ n }{ 2 }\left( a+l) \right)\] here n=7,a=26,l=44 calculate S7

OpenStudy (anonymous):

you just add them up the s7=245

OpenStudy (anonymous):

its more than just adding.. heres the example from the lesson: (1 + 100) + (2 + 99) + (3 + 98) + … + (49 + 52) + (50 + 51) = 50*101 = 5050.

OpenStudy (anonymous):

http://wn.com/infinite_arithmetic_series

OpenStudy (anonymous):

245.

OpenStudy (mathmale):

Angel Rose has come up with a cool resource: a video about arithmetic series. Some of you may be interested in another perspective: http://www.purplemath.com/modules/series4.htm This web page includes the formula suggested buy Surjithayer and provides examples as well.